OpenStudy (anonymous):

Determine all values of x, (if any), at which the graph of the function has a horizontal tangent. y(x)= (6x)/((x-9)^2)

OpenStudy (anonymous):

Help please

zepdrix (zepdrix):

Hey Becks c: So uhhh how far did you get? Looks like we need to apply quotient rule, yes?

zepdrix (zepdrix):

\[\Large\rm y(x)=\frac{6x}{(x-9)^2}\] \[\Large\rm y'(x)=\frac{\color{royalblue}{\left[6x\right]'}(x-9)^2-6x\color{royalblue}{\left[(x-9)^2\right]'}}{\left[(x-9)^2\right]^2}\]

zepdrix (zepdrix):

So there is our quotient rule setup, yah? :d

OpenStudy (anonymous):

yes, I got that far.

zepdrix (zepdrix):

Taking our derivatives, setting the y' equal to zero gives us,\[\Large\rm 0=\frac{\color{orangered}{\left[6\right]}(x-9)^2-6x\color{orangered}{\left[2(x-9)\right]}}{\left[(x-9)^2\right]^2}\]

zepdrix (zepdrix):

We want to look at the denominator and numerator `separately`.

zepdrix (zepdrix):

Recall that our original function had a \(\Large\rm (x-9)^2\) in the denominator. Which told us that x=9 is NOT in the domain of the function. So it turns out that the denominator of our derivative function isn't giving us any horizontal tangents. (Since 9 isn't a value we can consider). Actually there is another reason... critical values from the denominator usually correspond to vertical tangents, not horizontal.. blahhh let's not worry about that.

zepdrix (zepdrix):

Set your numerator equal to zero and solve for x.

zepdrix (zepdrix):

\[\Large\rm 0=6(x-9)^2-12x(x-9)\]

zepdrix (zepdrix):

What do you think Becks? Too confusing? c:

OpenStudy (anonymous):

Just a little bit. :p So the numerator doesn't have any horizontal tangents either?

zepdrix (zepdrix):

It doesn't? :o Hmm it probably does. Got to expand everything out in order to solve for x. Do you remember how to expand a square binomial?\[\Large\rm (x-9)^2=?\]

OpenStudy (anonymous):

(x-9)(x-9)

zepdrix (zepdrix):

Nooo silly, expand it out. FOIL it out.

OpenStudy (anonymous):

x^2-18x+81

zepdrix (zepdrix):

\[\Large\rm 0=6\color{green}{(x-9)^2}-12x(x-9)\]\[\Large\rm 0=6\color{green}{(x^2-18x+81)}-12x(x-9)\]Mmmm ok good.

zepdrix (zepdrix):

Distribute your 6, Distribute your -12x, and then combine like terms. What are you left with? Take a minute, try to get through those steps! c:

OpenStudy (anonymous):

9=x

zepdrix (zepdrix):

Well we already determined that x=9 isn't in the domain of our function. You're missing a value though :O If you simplified correctly, you should get:\[\Large\rm 0=486-6x^2\]Divide a 6 out,\[\Large\rm 0=81-x^2\]Which will factor to become,\[\Large\rm 0=(9-x)(9+x)\]

zepdrix (zepdrix):

If you had done it the other way:\[\Large\rm \frac{486}{6}=x^2\]Then you need to remember that when you take the `square root of a square`, you need a plop a +/- symbol in front.

OpenStudy (anonymous):

+/-9? so it would only be -9 because x= 9 is not in the domain of the function?

zepdrix (zepdrix):

Yessss, good :) x=-9 is our only solution for horizontal tangent line.

OpenStudy (anonymous):

thank you!

zepdrix (zepdrix):

yay team \c:/